OpenStudy (anonymous):

simplify : ( 3x^2 + 2xy - 1) - ( x^2 - 5x - y^2 + 6 ) HELPPPPPPPP....

OpenStudy (anonymous):

group it: (3x^2 -x^2)+2xy -(1+6) +5x+y^2=... can you take it from here?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

ok. let's try together... 2x^2 +2xy -7 +5x+y^2=(x+y)^2 +(x^2+5x-7) agreed so far?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

i'm not sure what grade you are in... what exactly you expected to do when "simplify"...? leave all as: 2x^2 +y^2+2xy +5x-7 or something else?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ok